While on an interplanetary mission, a 58.5-kg astronaut is floating toward the front of her ship at 0.15 m/s, relative to the ship. She wishes to stop moving, relative to the ship. She decides to throw away the 2.50-kg book she's carrying. What should the approximate speed and direction of the book be to achieve her goal?
Answer #1

Momentum conservation

In order to stop, she has to throw the book towards the front of ship

m1 = 58.5 kg

m2 = 2.5 kg

u1 = 0.15 kg

(m1+m2) u1 = m1 v1 + m2 v2

v1 = 0

So v2 = ( m1 + m2 ) u1 / m2

v2 = (58.5 + 2.5) x 0.15 / 2.5 = 3.66 m/s

==============================================================

Approximate speed = 3.66 m/s

Direction is towards the front of the ship